OpenStudy (anonymous):

Solve by factoring: y^2 + 2y = 8 can someone help me out here? D:

OpenStudy (anonymous):

very easy i will do

OpenStudy (anonymous):

bring all to one side: y^2+2y-8=0 factorise to get (y+4)(y-2)=0 that means y + 4 = 0, or y - 2 = 0 y = -4,2

OpenStudy (anonymous):

so then I would have to do y^2 + 2y - 8 = 8-8 and then y^2 + 2y - 8 = 0 how do I get it to (y + 4)(y - 2)=0?

OpenStudy (anonymous):

bring all to one side: y^2+2y-8=0 factorise to get (y+4)(y-2)=0 that means y + 4 = 0, or y - 2 = 0 y = -4,2

OpenStudy (anonymous):

y^2 + 2y - 8. We split the middle term. But first, what is the product of the first and last coefficients? (1 and -8) -8 Now, what two numbers multiply to get -8 and add to get 2 (the second coefficient) -2 and 4. y^2 - 2y + 4y - 8 = 0 Now group in pairs. y (y - 2) + 4 (y - 2) = 0 (y + 4) (y - 2) = 0